# OptiMind: Teaching LLMs to Think Like Optimization Experts

### v1

@@ A.1 Sample Problem Description and Output Optimization Model
Question: Maximize the total profit by determining the optimal production, inventory, and sales plan for three products (Product 0, Product 1, and Product 2) over a five-period planning horizon. The profit per unit sold is \$269 for Product 0, \$282 for Product 1, and \$241 for Product 2. The holding cost for each unit stored is \$15 per period. At the start of the planning horizon (Period 0), the production of each product must equal the sum of its sales and inventory for that period. For each subsequent period, the inventory carried over from the previous period plus the production in the current period must equal the sum of sales and inventory for the current period. The total production time required for all products on each machine type must not exceed the available capacity. The capacities are 480 hours per period for grinders, 320 hours per period for drills, and 160 hours per period for borers. Each unit of Product 0 requires 1 hour on grinders, 1 hour on drills, and 2 hours on borers. Each unit of Product 1 requires 1 hour on grinders, 1 hour on drills, and 2 hours on borers. Each unit of Product 2 requires 1 hour on grinders, 2 hours on drills, and 1 hour on borers. The inventory of each product at the end of any period cannot exceed 80 units. The maximum number of units that can be sold for each product in each period is as follows: Product 0: 48 in Period 0, 43 in Period 1, 58 in Period 2, 58 in Period 3, and 61 in Period 4; Product 1: 54 in Period 0, 56 in Period 1, 45 in Period 2, 46 in Period 3, and 40 in Period 4; Product 2: 57 in Period 0, 52 in Period 1, 68 in Period 2, 40 in Period 3, and 60 in Period 4. At the end of the planning horizon (Period 4), the inventory of each product must be exactly 20 units. Ensure that the final inventory levels meet the target requirements and that production does not exceed machine capacities. # Note: - The Code must include:“‘python import gurobipy as gp from gurobipy import GRB “‘ - Make sure the model variable is named ‘model‘. - Avoid using ”<<>>” in Gurobi constraints; instead, use ”<=<=” or ”>=>=” as appropriate. - Carefully determine whether the variable is an integer or a continuous variable.
@@
import gurobipy as gp
from gurobipy import GRB
@@
products = [0, 1, 2]
periods = [0, 1, 2, 3, 4]
machines = [’grinders’, drills’, borers’]
[…]
holding_cost = 15
[…]
grinders’: 480,
drills’: 320,
borers’: 160
[…]
grinders’: {0: 1, 1: 1, 2: 1},
drills’: {0: 1, 1: 1, 2: 2},
borers’: {0: 2, 1: 2, 2: 1}
[…]
0: {0: 48, 1: 43, 2: 58, 3: 58, 4: 61},
1: {0: 54, 1: 56, 2: 45, 3: 46, 4: 40},
2: {0: 57, 1: 52, 2: 68, 3: 40, 4: 60}
[…]
model = gp.Model(”Production_Planning”)
[…]
x = model.addVars(products, periods, name=”Production”)
s = model.addVars(products, periods, name=”Sales”)
I = model.addVars(products, periods, name=”Inventory”)
[…]
model.setObjective(
gp.quicksum(profit[i] * s[i, t] for i in products for t in periods) -
gp.quicksum(holding_cost * I[i, t] for i in products for t in periods),
GRB.MAXIMIZE
[…]
# Initial Inventory Balance (Period 0)
model.addConstrs((x[i, 0] == s[i, 0] + I[i, 0] for i in products), name=”Initial_Balance”)
[…]
model.addConstrs((I[i, t-1] + x[i, t] == s[i, t] + I[i, t] for i in products for t in periods if t >= 1), name=”Inventory_Balance”)
[…]
model.addConstrs((gp.quicksum(machine_time[m][i] * x[i, t] for i in products) <= machine_capacity[m] for m in machines for t in periods), name=”Machine_Capacity”)
[…]
model.addConstrs((s[i, t] <= max_sales[i][t] for i in products for t in periods), name=”Sales_Constraint”)
[…]
model.addConstrs((I[i, t] <= 80 for i in products for t in periods), name=”Inventory_Constraint”)
[…]
model.addConstrs((I[i, 4] == 20 for i in products), name=”Final_Inventory_Target”)
[…]
model.optimize()
[…]
if model.status == GRB.OPTIMAL:
print(”Optimal Solution Found!”)
print(fTotal Profit: \${model.ObjVal:.2f}”)
[…]
for t in periods:
print(fPeriod {t}: Production={x[i,t].X:.2f}, Sales={s[i,t].X:.2f}, Inventory={I[i,t].X:.2f}”)
else:
print(”No optimal solution found.”)
A.2 Problem Classification
@@
A.3 Prompts for Automatically Repairing Missing Data
@@ 3. The number of aircraft assigned does not exceed the availability of
A.4 Multi-Turn Inference Prompts
@@
A.5 Example of the Effect of Multi-Turn Inference
@@
A.6 Experiment Details
A.6.1 Training and Evaluation Setups
We performed Supervised fine-tuning (SFT) on a single node with eight NVIDIA HGX B200 GPUs. All solution generation in the data-cleaning pipeline and evaluation tasks were run on 4 compute nodes, each with eight 80GB NVIDIA H100 GPUs. We use gpt-oss-20b as the base model and adapt the Verl framework for SFT. Concretely, we train from the unsloth/openai-gpt-oss-20b-BFloat16 variant to avoid MXFP4 precision pitfalls and to leverage Unsloth’s compatibility fixes for gpt-oss, and uses FSDP2 strategy to distribute training across the eight GPUs. Key SFT hyperparameters and optimizer settings are listed in Table 7.
@@ A.6.1 Training and Evaluation Setups
For solution generation in data cleaning, we host openai/gpt-oss-20b with vLLM. For evaluation, we serve all SFT checkpoints and the BF16 base (unsloth/openai-gpt-oss-20b-BFloat16) via SGLang (vLLM did not support our Unsloth-adapted variant at the time when we conduct the evaluation). We use top-p decoding for all generation and inference; the common sampling settings are summarized in Table 8.
@@
A.7 Details of Error-Analysis Hints at Inference
@@
A.8 Details of Training Data Cleaning
@@
A.9 Details of Benchmarks Cleaning
This section dives into some of the issues we observed in the benchmarks data. These fall under six categories: missing data, ambiguous problem descriptions, integral vs fractional variables, wrong solutions, infeasible problems, and out-of-scope problems. We next provide additional details specific examples that highlight these challenges.
A.9.1 Missing Data
@@
A.9.2 Ambiguous Problem Descriptions
@@
A.9.3 Integral vs Fractional Variables
Many problems exhibit ambiguity regarding whether decision variables are integer or continuous. For example, a production optimization problem in IndustryOR refers to “units of products”, while the reported ground truth corresponds to the production of fractional products. We resolve this by either updating the ground truth value, adding explicit clarification sentences that enforce integrality, or when both interpretations are reasonable, by providing two ground-truth answers corresponding to the integer and fractional formulations.
@@ A.9.5 Infeasible Problems
A.9.6 Out-of-Scope Problems
Finally, we also observe a small number of non-linear problems. For instance, OptMATH contains certain second-order cone programming and quadratically constrained programming problems. Since our focus in this work is on (mixed integer) linear optimization, we manually omit these non-linear instances from the benchmarks. We also find that the available training datasets contain no comparable nonlinear training data, so including such instances would create a train–test mismatch.

### v2

@@ A.2 Sample Problem Description and Output Optimization Model
Question: Maximize the total profit by determining the optimal production, inventory, and sales plan for three products (Product 0, Product 1, and Product 2) over a five-period planning horizon. The profit per unit sold is \$269 for Product 0, \$282 for Product 1, and \$241 for Product 2. The holding cost for each unit stored is \$15 per period. At the start of the planning horizon (Period 0), the production of each product must equal the sum of its sales and inventory for that period. For each subsequent period, the inventory carried over from the previous period plus the production in the current period must equal the sum of sales and inventory for the current period. The total production time required for all products on each machine type must not exceed the available capacity. The capacities are 480 hours per period for grinders, 320 hours per period for drills, and 160 hours per period for borers. Each unit of Product 0 requires 1 hour on grinders, 1 hour on drills, and 2 hours on borers. Each unit of Product 1 requires 1 hour on grinders, 1 hour on drills, and 2 hours on borers. Each unit of Product 2 requires 1 hour on grinders, 2 hours on drills, and 1 hour on borers. The inventory of each product at the end of any period cannot exceed 80 units. The maximum number of units that can be sold for each product in each period is as follows: Product 0: 48 in Period 0, 43 in Period 1, 58 in Period 2, 58 in Period 3, and 61 in Period 4; Product 1: 54 in Period 0, 56 in Period 1, 45 in Period 2, 46 in Period 3, and 40 in Period 4; Product 2: 57 in Period 0, 52 in Period 1, 68 in Period 2, 40 in Period 3, and 60 in Period 4. At the end of the planning horizon (Period 4), the inventory of each product must be exactly 20 units. Ensure that the final inventory levels meet the target requirements and that production does not exceed machine capacities. # Note: - The Code must include:“‘python import gurobipy as gp from gurobipy import GRB “‘ - Make sure the model variable is named ‘model‘. - Avoid using "<<" >>" in Gurobi constraints; instead, use "<=<=" or ">=>=" as appropriate. - Carefully determine whether the variable is an integer or a continuous variable.
@@
1import gurobipy as gp
2from gurobipy import GRB
3
4# Data
5products = [0, 1, 2]
6periods = [0, 1, 2, 3, 4]
7machines = ['grinders', 'drills', 'borers']
8
9profit = {0: 269, 1: 282, 2: 241}
10holding_cost = 15
11
12machine_capacity = {
13 'grinders': 480,
14 'drills': 320,
15 'borers': 160
16}
17
18machine_time = {
19 'grinders': {0: 1, 1: 1, 2: 1},
20 'drills': {0: 1, 1: 1, 2: 2},
21 'borers': {0: 2, 1: 2, 2: 1}
22}
23
24max_sales = {
25 0: {0: 48, 1: 43, 2: 58, 3: 58, 4: 61},
26 1: {0: 54, 1: 56, 2: 45, 3: 46, 4: 40},
27 2: {0: 57, 1: 52, 2: 68, 3: 40, 4: 60}
28}
29
30# Model
31model = gp.Model("Production_Planning")
32
33# Decision Variables
34x = model.addVars(products, periods, name="Production")
35s = model.addVars(products, periods, name="Sales")
36I = model.addVars(products, periods, name="Inventory")
37
38# Objective Function
39model.setObjective(
40 gp.quicksum(profit[i] * s[i, t] for i in products for t in periods) -
41 gp.quicksum(holding_cost * I[i, t] for i in products for t in periods),
42 GRB.MAXIMIZE
43)
44
45# Constraints
46# Initial Inventory Balance (Period 0)
47model.addConstrs((x[i, 0] == s[i, 0] + I[i, 0] for i in products), name="Initial_Balance")
48
49# Inventory Balance (Periods 1-4)
50model.addConstrs((I[i, t-1] + x[i, t] == s[i, t] + I[i, t] for i in products for t in periods if t >= 1), name="Inventory_Balance")
51
52# Machine Capacity Constraints
53model.addConstrs((gp.quicksum(machine_time[m][i] * x[i, t] for i in products) <= machine_capacity[m] for m in machines for t in periods), name="Machine_Capacity")
54
55# Sales Constraints
56model.addConstrs((s[i, t] <= max_sales[i][t] for i in products for t in periods), name="Sales_Constraint")
57
58# Inventory Constraints
59model.addConstrs((I[i, t] <= 80 for i in products for t in periods), name="Inventory_Constraint")
60
61# Final Inventory Target
62model.addConstrs((I[i, 4] == 20 for i in products), name="Final_Inventory_Target")
63
64# Optimize
65model.optimize()
66
67# Results Interpretation
68if model.status == GRB.OPTIMAL:
69 print("Optimal Solution Found!")
70 print(f"Total Profit: \${model.ObjVal:.2f}")
71
72 for i in products:
@@
74 for t in periods:
75 print(f"Period {t}: Production={x[i,t].X:.2f}, Sales={s[i,t].X:.2f}, Inventory={I[i,t].X:.2f}")
76else:
77 print("No optimal solution found.")
A.3 Problem Classification
@@
A.4 Prompts for Automatically Repairing Missing Data
@@ 3. The number of aircraft assigned does not exceed the availability of
A.5 Multi-Turn Inference Prompts
@@
A.6 Example of the Effect of Multi-Turn Inference
@@
A.7 Experiment Details
We performed Supervised fine-tuning (SFT) on a single node with eight NVIDIA HGX B200 GPUs. All solution generation in the data-cleaning pipeline and evaluation tasks were run on 4 compute nodes, each with eight 80GB NVIDIA H100 GPUs. We use gpt-oss-20b as the base model and adapt the Verl framework for SFT. Concretely, we train from the unsloth/openai-gpt-oss-20b-BFloat16 variant to avoid MXFP4 precision pitfalls and to leverage Unsloth’s compatibility fixes for gpt-oss, and uses FSDP2 strategy to distribute training across the eight GPUs. We tune the learning rate over $10^{-6},4\times 10^{-5}]$ via a grid search, and select the best configuration on the validation split under the single-turn, no-hint setting. Key SFT hyperparameters and optimizer settings are listed in Table 6.
@@ A.7 Experiment Details
For solution generation in data cleaning, we host openai/gpt-oss-20b with vLLM. For evaluation, we serve all SFT checkpoints and the BF16 base (unsloth/openai-gpt-oss-20b-BFloat16) via SGLang, since vLLM did not support our Unsloth-adapted variant when we conducted the evaluation. We use top-p decoding for all generation and inference. We tune the sampling temperature over $[0.5,1.0]$ via a grid search, and select the best configuration on the validation split under the single-turn, no-hint setting. The common sampling settings for our SFT model are summarized in Table 7.
@@
A.8 Details of Error-Analysis Hints at Inference
@@
A.9 Details of Training Data Cleaning
@@
A.10 Details of Benchmarks Cleaning
This section describes representative issues we observed in the benchmark data. These fall into six main categories: missing data, ambiguous problem descriptions, integral vs. fractional variables, wrong solutions, infeasible problems, and out-of-scope problems. We next provide additional details and examples that highlight these challenges.
[…]
Table 8 shows representative examples of incorrect problems and their fixes in IndustryOR and OptMATH, while Table 9 summarizes the frequency and distribution of major error types across both benchmarks. A complete, per-instance correction record is available in the HTML summary tables described in our supplementary material.
Table 8: Representative incorrect problems and fixes in IndustryOR and OptMATH.
Benchmark Problem Index Optimization Class Why the original problem is incorrect How we fixed it
IndustryOR 24 The problem did not state an objective function. We added a well-specified objective (minimize pollution).
IndustryOR 46 The original instance was infeasible but had a ground-truth objective value of $472.3$. We updated the data to make the instance feasible and adjusted the ground-truth objective accordingly.
OptMATH 35 Facility Location Problem (FLP) The problem describes selecting towers to cover regions, but the description is missing data about which towers cover which regions. We generated the missing coverage data semi-automatically and then verified it manually.
OptMATH 36 Job Shop The instance omits the number of machines and specifies only a partial assignment of jobs to machines. We generated the missing machine and assignment data semi-automatically and then verified it manually.
Table 9: High-level breakdown of benchmark errors. Some instances belong to multiple error categories. Percentages are relative to the total number of problems in each benchmark.
Benchmark # (%) Incorrect Missing Data Ambiguous Infeasible Wrong GT Underspecified Specific Solver Non-linear Inconsistent Integral / Fractional
IndustryOR 50 (50%) 0 (0%) 11 (11%) 2 (2%) 25 (25%) 5 (5%) 5 (5%) 1 (1%) 2 (2%) 4 (4%)
OptMATH 68 (41%) 48 (29%) 11 (7%) 1 (0.6%) 10 (6%) 1 (0.6%) 0 (0%) 0 (0%) 0 (0%) 0 (0%)
Cleaning these benchmarks is nontrivial given the high initial error rate in the original data sources. To ensure rigor, we conduct a total of four passes of expert review. All corrections are documented in the supplementary material, and we plan to release the full cleaning logs on GitHub for broader visibility.
A.10.1 Missing Data
@@
A.10.2 Ambiguous Problem Descriptions
@@
A.10.3 Integral vs Fractional Variables
Many problems exhibit ambiguity regarding whether decision variables are integer or continuous. For example, a production optimization problem in IndustryOR refers to “units of products”, while the reported ground truth corresponds to the production of fractional products. We resolve this by updating the ground truth value, adding explicit clarification sentences that enforce integrality or continuity.
@@ A.10.5 Infeasible Problems
A.10.6 Out-of-Scope Problems
Finally, we also observe a small fraction of non-linear problems. For instance, OptMATH contains certain second-order cone programming and quadratically constrained programming problems. We deliberately omit these non-linear instances from the benchmarks for the following reasons:
First, all second-order test problems in the benchmark are essentially copies of the same synthetic template (see the example below). In addition, these instances suffer from severe missing-data issues: many problem statements mention, for example, the OptMATH problems with index 142, 148, 150 and 154 mentioned “12 linear constraints” but only specify parameter values for one or two constraints, so the problems are incomplete and require substantial manual reconstruction by human labelers before they can be meaningfully used.
Second, when we do run our models on these nonlinear instances, we find that the errors are very concentrated: the solver typically encodes $\|y_{k}\|_{2}\leq t_{k}^{2}$ constraints with $t_{k}$ unrestricted, making the feasible region a nonconvex union of two cones. This is essentially modeling issue that we believe would be straightforward to fix if similar second-order training examples (with correctly encoded SOCs and linking equations) were available. However, **we do not find any comparable second-order problems of this type in our training data**. This means we cannot derive targeted, class-specific hints from the training corpus to address these errors, in contrast to the MILP classes where our method is designed to operate.
The closest problem family we see in the available datasets is classical mean–variance portfolio optimization with a quadratic objective and linear constraints, e.g.
$\min_{w}$ $\sum_{i=1}^{n}\sum_{j=1}^{n}w_{i}w_{j}\sigma_{ij}\$s.t. $\sum_{i=1}^{n}w_{i}r_{i}\geq R,$ $\sum_{i=1}^{n}w_{i}=1,\$$l_{i}\leq w_{i}\leq u_{i}\ \forall i.$
whose description and structure are quite different from the synthetic second-order cone instances in the test set. In practice, hints or error patterns derived from such portfolio problems do not transfer to the benchmark’s second-order templates.
Because we build upon publicly available datasets in which second-order problems are extremely scarce and not representative of the nonlinear test instances, we view a thorough treatment of these problems as out of scope for this work. That said, our framework is general: given a sufficiently rich set of similar nonlinear problems in the training data, we expect that the same error-analysis-and-hints pipeline could be applied to derive effective hints and substantially reduce these errors as well.
@@ A.10.6 Out-of-Scope Problems
Supplementary Material: Released Cleaned Test Sets
We release the expert–cleaned test sets used in our evaluation as supplementary material:
  • optimind_cleaned_industryor.csv (99 items),
  • optimind_cleaned_mamo_complex.csv (210 items), and
  • optimind_cleaned_optmath.csv (130 items),
Each CSV is UTF–8 encoded and contains two columns: question (the complete natural–language problem statement) and answer (the ground–truth optimal objective value after cleaning).
The answer field is either (i) a single numeric value, or (ii) a JSON array of numeric values when multiple interpretations are reasonable (e.g., integer vs. fractional formulations) and thus multiple objectives are accepted. During evaluation, a prediction is considered correct if it matches any provided value within absolute and relative tolerances of $10^{-6}$.
These files reflect our corrections for missing parameters, ambiguity, infeasibility, wrong reference answers, and scope mismatches identified during manual review. The sets are intended strictly for evaluation (not training) to ensure comparability across studies. We will update the archive if community feedback identifies additional issues; any changes will be documented with version notes.
Supplementary Material: Dataset Correction Tables
In addition to the cleaned CSV test sets, for the IndustryOR and OptMATH benchmarks we provide HTML summary tables that list each correction and the reason for the change. Each HTML file contains a table with columns [Problem Index, Original Problem, Original Answer, Updated Problem, Updated Answer, How did we fix it]. Figure 14 shows a screenshot of one such table, illustrating the layout and content of the HTML files. We also release comparison files industryOR_original_vs_ours.html and OptMATH_original_vs_ours.html, which align the original instances with our cleaned versions. In addition, compare_SIRL_Ours.html compares our cleaned IndustryOR set against the SIRL-cleaned version: although the latter was reported as corrected, we identify five instances that still exhibit issues and document them in this file. Together, these tables offer a clear, human-readable view of the structure and details of our dataset corrections.
Figure 14: A screenshot of the HTML correction table in our supplementary material showing our edits for the IndustryOR benchmark.